# Premerger phenomena in neutron-star binary coalescences

### v2

@@ V.5 Waiting times
It is evident that there is a significant spread of precursor phenomena, with a couple orders of magnitude separating the shortest and longest waiting times and also durations. It is probable therefore that there are multiple ignition mechanisms, involving either pre- or post-merger phenomena; the physics pertaining to delays are discussed in Secs. VI.1 and VI.3.3. We remark on the nature of some entries in Tab. 3, as some have error bars while others do not. In Wang et al. 425, from which many of our tabulated precursor results are quoted from, the authors did not given a confidence interval attached to the $\pm$ values they provided. As per standard convention, these error bars may correspond to $84.13\%$ confidence upper and lower limits containing the $68.27\%$ confidence interval [155]. In some other studies, such as Minaev and Pozanenko 276, Zhong et al. 448, waiting times are given as a kind of mean value from $T_{90}$ data (see Sec. V.1), which we quote here using the $\sim$ symbol. Owing to these differences in notation and convention, on top of other systematics described thus far, caution should be applied when weighting precursor significance across different studies.
Table 3: Precursors (likely) associated with merger events, organised in descending order of waiting timing relative to main event, with remarks on special features (if any). Key: XAM = X-ray afterglow with internal plateau, suggestive of a magnetar remnant as identified by Rowlinson et al. 340 and others [e.g. 384]. Most data were compiled by Wang and Liu 422, though others not listed there come from Minaev and Pozanenko 276 (071030), Zhong et al. 448 (100717, 130310A, 170726794), Dichiara et al. 119 (230307A), and Deng et al. 112 (180703B). The long-form Fermi GBM catalogue name is used for some GRBs. Some precursors may be postmerger.
@@
071030 $0.9\pm 0.2$ $\sim 2.5$ $2.7\pm 0.5$ Data appear lost [see 448]? Debated $T_{90}$
@@ V.6.1 GRB 211211A
GRB 211211A was accompanied by a kilonova [328]. This by itself is exciting: the GRB was long $(T_{90}\gtrsim 30$s), and yet the association with a kilonova identifies the origin as a merger event (see Sec. V.2). Aside from this, the event showed a precursor flare $\sim 1$s prior to the main event [433]. In Suvorov et al. 385, the precursor timing was matched with a $g$-mode resonance with sufficiently large overlap that the crust may break, within the resonant failure model (see Sec. VI.3). Since magnetic fields skew the mode eigenfrequency, the magnetic field strength that was obtained $(\gtrsim 10^{14}$ G) was required to be consistent with that set by the extraordinarily high isotropic luminosity $(\sim 7\times 10^{49}\text{erg s}^{-1}$) and the absence of thermal emissions in the precursor (see Sec. V.4). Aside from this, the event was modulated by quasi-periodic oscillations (QPOs) at $\sim 22$ Hz [433] [see also 87].
Seismic excitations in the crust is a natural prediction of a failure: a deep quake relieves energy from the crust which then causes the surroundings to vibrate through aftershocks. These aftershocks are most conducive to torsional modes, as initially pointed out by Duncan 130, which happen to have frequencies in the $\sim 20$ Hz range. The combination of having mass estimates for the binary members from the kilonova [328], estimates for the internal properties and magnetic field strength from a $g$-mode resonance, and estimates for the shear modulus in the crust from the torsional mode frequency proves intriguing. Although a thorough parameter space scan was not performed by Suvorov et al. 385, it was shown that conventional assumptions on the EOS (Fig. 2) and shear modulus from recent literature can accommodate all of the observed properties. More precisely, the frequency of the precursor QPO (interpreted as a torsional mode) and relative timing of the precursor itself (interpreted as a $g$-mode resonance) can be well-fit with a star that has kilonova-inferred mass $M=1.25M_{\odot}$ [328] with the APR EOS (Sec. II.1.2) and shear modulus computed from nuclear physics (Sec. III.1). Other interpretations are valid, though. For example, the QPO could be associated with a magnetoelastic mode [e.g. 150, 151] and the precursor could be attributed to an interface-mode [e.g. 407, 406, see also Sec. VI.4.1].
@@ V.6.5 GRB 090510
The two precursors from this event were separated by $\sim 12$ s [403], with the second precursor taking place $\sim 0.5$ s prior to the main event. Several possibilities present themselves, involving (i) two premerger precursors, (ii) two postmerger precursors, or (iii) one of each. The parameter space is large. The first precursor was notably softer ($\sim 30$ keV peak) than the second ($\sim 300$ keV peak), which disfavours the operation of strong external magnetic fields at early stages. A possibility put forward by Kuan et al. 232 involves rotation and stellar misalignment. As shown in Figure 15, a significant misalignment angle and rapid rotation (though the latter is not strictly necessary, even spins of $\lesssim 10$ Hz are sufficient) can instigate two episodes of crustal failure with $m=2$ and $m=1$ modes (see Sec. IV.3). Although the delay seen in this Figure does not exactly correspond to that seen in 090510, it demonstrates the possibility. Regarding the 090510 double event specifically, parameter-space constraints are given in Kuan et al. 232.
Figure 15: A double resonance of $m=2$ and $m=1$ kink modes in a highly misaligned and rapidly-rotating neutron star. The strain ($\sigma$) can exceed $\sim 0.04$ — a realistic elastic maximum [42, 222, 41] — and thus, theoretically, trigger two episodes of failure separated by a number of seconds. The stratification rises to $\delta\approx 0.021$ for the later episode, due to intervening tidal heating (cf. Sec. IV.5), having started from an initial (cold) value of $\delta\approx 0.006$. The binary is symmetric with stars of mass $1.23M_{\odot}$ and the APR4 EOS. From Kuan et al. 232.
The second remarkable feature of GRB 090510 is that of a rather prominent X-ray afterglow, as discussed by Rowlinson et al. 340, Suvorov and Kokkotas 382 and others. This afterglow exhibited the ‘internal plateau’ feature thought to be characteristic of a neutron star pumping spindown energy into the forward shock. Importantly, the system showed a steady decline in flux rather than a sharp cutoff, which is often interpreted as the collapse time of the metastable star [see, e.g. 330, 379]. If indeed a stable neutron star was formed, it is more plausible that the jet formation timescale was comparatively long, and hence that only one precursor was premerger with the second being attributable to a choking jet, possibly because neutron-stars have a harder time forming jets with high Lorentz factors [93]. Still, several scenarios are viable.
@@ V.6.6 GRB 230307A
GRB 230307A is an extremely bright, long GRB. The recorded gamma-ray fluence is $\gtrsim 3\times 10^{-3}\text{ erg cm}^{-2}$ in the 10–1000 keV band [119], second only to GRB 221009A. Despite its long duration, it is likely associated with a kilonova and heavy element production [244], thus resembling the case of GRB 211211A [119] of a merger-driven long GRB. Taking the best-fit distance of $\approx 291$ Mpc implies an enormous luminosity $L_{\rm prec}\sim 3.6\times 10^{50}\text{ erg s}^{-1}$. As discussed by Dichiara et al. 119, this event is a strong candidate for a magnetar with an external field of $\gtrsim 10^{15}$ G taking place in a merger. If the field was of this magnitude, accounting for magnetic shifting in mode eigenfrequencies would be essential, as would electromotive torques towards the end of the binary evolution (see Sec. IV.4). Given the waiting time of only $\approx 0.4$ s however, many ignition scenarios are viable (especially post-merger ones).
@@ V.7 Connections with post-merger phenomena
Insofar as strong magnetic fields are necessary to explain premerger, precursor spectra (Sec. V.4), the remnant environment and structure could be affected. It is fair to say that one of the open problems in GRB physics concerns whether or not magnetar remnants are capable of launching a relativistic jet that can drill through the polar baryon pollution [e.g. 345]. Most numerical simulations of mergers that leave a magnetar find that (i) a magnetic field of strength approaching equipartition ($\sim 10^{17}$ G) is needed to a launch a jet (if a jet can be launched at all), and (ii) the magnetic field that is generated is larger if the pre-merger ‘seed’ fields are larger [e.g. 92]. Point (ii) has been challenged by recent, high-resolution simulations [212], though it is naturally easier to accept strong remnant fields via flux conservation arguments if there is a magnetar in the merger.
The magnetic field strength of a remnant star would also influence its spindown energy and GW luminosity. In principle, one could anticipate a subpopulation of remnants with stronger fields if bright precursors are observed therefore [386]. Even if irrelevant for precursors, dynamo activity triggered by $f$- or other mode resonances just before merger could also amplify seed fields (see Sec. VI.5).
@@ VI.1 Delay Timescales: postmerger jets
Fortunately — at least from a premerger perspective — there is reason to suspect that $t_{\rm J}$ may be considerably shorter than that seen in GRB 170817A in some instances, perhaps even down to a $\sim$ few ms. In this section we closely follow and summarise the compilations of Zhang 446 and Burns 74 in describing various jet formation mechanisms and $t_{\rm J}$ estimates. Where appropriate, some recent numerical simulations or theoretical results concerning remnants are instead quoted to update and/or adjust values presented in Table 1 of Zhang 446.
@@
Zhang 446 introduces a second decomposition for $\Delta t_{\rm jet}$, depending on the mechanism responsible for launching. The mechanism is naturally dependent on the progenitors. We write
@@
Table 4: Expected delays associated with merger-driven GRBs and relative precursor timings, adapted and expanded from Zhang 446. Abbreviations: NS = Neutron star; SM = Supramassive; HM = hypermassive. Remarks: (*) HMNS collapse time could theoretically be prolonged through thermal [300] or magnetic [379] support, extending the upper limit significantly. (**) Result quoted from analysis performed in Ciolfi et al. 93 regarding breakout requirements [see also Mösta et al. 281 and Ciolfi 91 for detailed discussions]. (†) Results anticipated from high-resolution dynamo simulations of Kiuchi et al. 212, who found amplification on sub-ms timescales. (‡) From simulations of Ruiz et al. 343, who found that sufficiently low-mass BHNSs result in a HMNS that tends to launch a magnetically-dominated jet. Argued by Shibata and Taniguchi 355 that for canonical neutron-star parameters, SGRB is limited to “low” energies of order $\sim 10^{48}$ erg/s because of tidal disruption impact on torus mass. (#) We assume a fireball-like mechanism for accretion scenarios involving a neutron star, or a Blandford-Znajek mechanism if involving a BH, and a Poynting-flux-dominated jet for dynamo (magnetic) scenarios when estimating $\Delta t_{\rm GRB}$. Though, in principle, any of these mechanisms could have an upper limit of many $\sim$ s owing to theoretical uncertainties; see Zhang 446 and Burns 74. Black hole topology and modified-gravity terms could also skew these estimates, though such effects are ignored [see, e.g. 362, 284].
@@
We close this section by noting that multimessenger events may accompany a collapse event for some of the remnant types listed in Tab. 4. Most notably, fast radio bursts (FRBs) have been predicted to be produced following the collapse of a neutron star [see, e.g. 136, 379]. Once an event horizon forms and cloaks the star, the external field lines are no longer anchored and so snap, inciting relativistic, magnetic shocks that accelerate electrons to high Lorentz factors, ultimately producing radiation in the $\gtrsim$ GHz band. A coincident detection of a sharp drop in X-ray flux and an FRB would be a “smoking gun” for a post-merger, neutron-star collapse; in such a case, at least $\Delta t_{\rm rem}$ could be tightly constrained. See Lu et al. 251 for recent (unsuccessful) search efforts for coincident FRBs and GRBs, though see Moroianu et al. 277 who found that FRB 20190425A is coincident with GW190425 at $\sim 2.8\sigma$. Sadly, no GRB was observed for this event.
@@ VI.2 Magnetospheric interaction and unipolar inductor
Depending on the persistence (or late-time generation) of strong (crustal) magnetic fields over cosmological timescales, electrodynamic interactions occurring in the final seconds of inspiral could produce fireworks. The mechanisms responsible have been described in detail by Hansen and Lyutikov 177 and others since. The production of gamma-rays can proceed through essentially three different channels, depending on the relative magnetization and alignment angles of the two stars [312, 235, 424]. In cases with a black hole, comparable estimates can be found in McWilliams and Levin 264.
The simplest, and arguably most probable if a luminous precursor is observed, scenario in the electrodynamic interaction context is that of the unipolar inductor introduced in Sec. IV.4. The reason for this is that it is already difficult to explain one magnetar in a merger let alone two, and without strong fields to mediate the interaction the precursor is likely to fizzle out long before it reaches Earth. The other two cases instead apply when the object’s have comparable magnetic dipole moments, either aligned or anti-aligned (though in reality of course there will be some angle, $\Theta_{\mu}$). In the aligned case, the field lines will compress at the interaction radius given by $r_{i}\approx a/(1+q_{\mu}^{1/3})$ with dipole-ratio $q_{\mu}=\mu_{B,d}/\mu_{A,d}$ [424]. In the opposite case, we have instead direct reconnection occurring in the interaction zone which leads to explosive event(s) followed by the establishment of a quasi-stable circuit similar to a unipolar inductor [e.g. 293, 100].
Lai 235 argues that the maximum energy that can be relieved relates to $\dot{E}_{\rm EM}$ from equation (47). This is controlled by the effective resistance between the two stars through the effective twist parameter $\zeta_{\phi}$, viz.
@@
where $B$ here is to be understood as the largest equatorial strength of either star and we have assumed equal stellar radii. The above is difficult to reconcile with precursor luminosities unless $a\sim 3R$ and $\zeta_{\phi}\gg 1$, though $\dot{E}_{\rm EM}$ can be slightly larger than estimates provided by Wang et al. 424 if $\zeta_{\phi}$ is kept free.
The estimate (55) is in rough agreement with that found from ab initio numerical simulations for large twists. Most and Philippov 279 found, via special-relativistic simulations, a maximum dissipation of (equation 6 therein)
@@
applying just $\sim$ ms before merger for equal mass objects [see also 100]. Note in particular the relevant $B$ value here is the local value at the interaction radius not that at the stellar surface (though these are approximately equal at merger by definition). Interestingly, the luminosity scaling was found by Most and Philippov 279 to be $\propto a^{-7/2}$, which is shallower than that of (55). Either way, we can match (56) with (55) at merger if $\zeta_{\phi}\lesssim 10$, consistent with the anti-aligned simulations by Palenzuela et al. 293, Crinquand et al. 100. Most and Philippov 280, Carrasco et al. 76 studied neutron-star plus black-hole binaries, finding roughly similar dissipation rates though with the interesting result that the flares present in their simulations were up to $\sim 10^{2}$ times brighter than the (relativistic) orbital emission.
An important aspect of the interaction model is that emissions should be generated almost instantaneously. That is, particle production occurs immediately after reconnection and thus the precursor timing should correspond directly to the ignition moment without an additional lag beyond those detailed in Sec. VI.1. On the other hand, the duration of a precursor born from magnetospheric interaction is likely limited to $\sim 10$ ms based on the relationship between the Poynting flux and chirp length where the separations are low [139]. This is somewhat shorter than all event durations listed in Tab. 3, though allowing for a factor $\sim$ few extension would bring many low-waiting-time precursors into view (e.g. 150922A, 100223110 , 100702A, …). Still, Metzger and Zivancev 272 argue the detection horizon is limited to $\sim 10$ Mpc for $B\approx 10^{14}$ G unless the twist is large, $\zeta_{\phi}\gg 1$.
Finally, aside from reconnection being generally able to spark a plethora of progenitors (see also Sec. VII), strong magnetic fields in the emission environment could effect photon propagation directly. For example, quantum-mechanical photon splitting could acquire a non-negligible cross section [109, 265], thereby impacting on the polarization states and/or Faraday rotation of electromagnetic waves. Such effects, manifesting as vacuum birefringence, have been observed in emissions from the isolated neutron stars RX J1856.5-3754 [275] and 4U 0142+61 [396] with characteristic field strengths of a few by $10^{13}$ G [317] and $\sim 10^{14}$ G [291], respectively. As noted by Wang and Liu 422, the next-generation gamma-ray polarimeter POLAR-2 may be able to measure such effects directly and constrain the plasma environment [195].

### v3

@@ V.5 Waiting times
It is evident that there is a significant spread of precursor phenomena, with a couple orders of magnitude separating the shortest and longest waiting times and also durations. It is probable therefore that there are multiple ignition mechanisms, involving either pre- or post-merger phenomena; the physics pertaining to delays are discussed in Secs. VI.1 and VI.3.3. We remark on the nature of some entries in Tab. 3, as some have error bars while others do not. In Wang et al. 439, from which many of our tabulated precursor results are quoted from, the authors did not given a confidence interval attached to the $\pm$ values they provided. As per standard convention, these error bars may correspond to $84.13\%$ confidence upper and lower limits containing the $68.27\%$ confidence interval [159]. In some other studies, such as Minaev and Pozanenko 287, Zhong et al. 463, waiting times are given as a kind of mean value from $T_{90}$ data (see Sec. V.1), which we quote here using the $\sim$ symbol. Owing to these differences in notation and convention, on top of other systematics described thus far, caution should be applied when weighting precursor significance across different studies.
Table 3: List of precursors (likely) associated with merger events, organised in descending order of waiting timing relative to main event, with some remarks on special features (if any). Key: XAM = X-ray afterglow with internal plateau, possibly suggestive of a magnetar remnant as identified by Rowlinson et al. 354 and others [e.g. 396, 398, 399]. Most data were compiled by Wang and Liu 436, though others not listed there come from Minaev and Pozanenko 287 (071030), Zhong et al. 463 (100717, 130310A, 170726794), Dichiara et al. 119 (230307A), and Deng et al. 112 (180703B). The long-form Fermi GBM catalogue name is used for some GRBs. Some precursors may be postmerger.
@@
071030 $0.9\pm 0.2$ $\sim 2.5$ $2.7\pm 0.5$ Data appear lost [see 463]? Debated $T_{90}$
@@ V.6.1 GRB 211211A
GRB 211211A was accompanied by a kilonova [341]. This by itself is exciting: the GRB was long $(T_{90}\gtrsim 30$s), and yet the association with a kilonova identifies the origin as a merger event (see Sec. V.2). Aside from this, the event showed a precursor flare $\sim 1$s prior to the main event [447]. In Suvorov et al. [399], the precursor timing was matched with a $g$-mode resonance with sufficiently large overlap that the crust may break, within the resonant failure model (see Sec. VI.3). Since magnetic fields skew the mode eigenfrequency, the magnetic field strength that was obtained $(\gtrsim 10^{14}$ G) was required to be consistent with that set by the extraordinarily high isotropic luminosity $(\sim 7\times 10^{49}\text{erg s}^{-1}$) and the absence of thermal emissions in the precursor (see Sec. V.4).
In addition to that described thus far, the duration of the main event may also point towards magnetar participation in the merger. Kiuchi et al. [220] found in GRMHD simulations with merging magnetars that the accretion disc that comes to surround the remnant is itself highly magnetised. Although not seen in their simulation, it has been proposed that this, combined with the magnetic field sourced by the remnant if a neutron star, can set up a barrier that episodically chokes the accretion flow thereby prolonging the event duration [see 447, and references therein].
The 211211A precursor was also modulated by quasi-periodic oscillations (QPOs) at $\sim 22$ Hz [447] [see also 86]. Seismic excitations in the crust is a natural prediction of a failure: a deep quake relieves energy from the crust which then causes the surroundings to vibrate through aftershocks. These aftershocks are most conducive to torsional modes, as initially pointed out by [132], which happen to have frequencies in the $\sim 20$ Hz range. The combination of having mass estimates for the binary members from the kilonova [341], estimates for the internal properties and magnetic field strength from a $g$-mode (or other) resonance, and estimates for the shear modulus in the crust from the torsional mode frequency proves intriguing. Although a thorough parameter space scan was not performed by Suvorov et al. [399], it was shown that conventional assumptions on the EOS (Fig. 2) and shear modulus from recent literature can accommodate all of the observed properties. More precisely, the frequency of the precursor QPO (interpreted as a torsional mode) and relative timing of the precursor itself (interpreted as a $g$-mode resonance) can be well-fit with a star that has kilonova-inferred mass $M=1.25M_{\odot}$ [341] with the APR EOS (Sec. II.1.2) and shear modulus computed from nuclear physics (Sec. III.1). Other interpretations are valid, though. For example, the QPO could be associated with a magnetoelastic mode [e.g. 153, 154] and the precursor could be attributed to an interface-mode [e.g. 421, 420, see also Sec. VI.4.1].
@@ V.6.5 GRB 090510
The two precursors from this event were separated by $\sim 12$ s [417], with the second precursor taking place $\sim 0.5$ s prior to the main event. Several possibilities present themselves, involving (i) two premerger precursors, (ii) two postmerger precursors, or (iii) one of each. The parameter space is large. The first precursor was notably softer ($\sim 30$ keV peak) than the second ($\sim 300$ keV peak), which disfavours the operation of strong external magnetic fields at early stages. A possibility put forward by [241] involves rotation and stellar misalignment. As shown in Figure 15, a significant misalignment angle and rapid rotation (though the latter is not strictly necessary, even spins of $\lesssim 10$ Hz are sufficient) can instigate two episodes of crustal failure with $m=2$ and $m=1$ modes (see Sec. IV.3). Although the delay seen in this Figure does not exactly correspond to that seen in 090510, it demonstrates the possibility. Regarding the 090510 double event specifically, parameter-space constraints are given in [241].
Figure 15: A double resonance of $m=2$ and $m=1$ kink modes in a highly misaligned and rapidly-rotating neutron star. The strain ($\sigma$) can exceed $\sim 0.04$ — a realistic elastic maximum [42, 230, 41] — and thus, theoretically, trigger two episodes of failure separated by a number of seconds. The stratification rises to $\delta\approx 0.021$ for the later episode, due to intervening tidal heating (cf. Sec. IV.5), having started from an initial (cold) value of $\delta\approx 0.006$. The binary is symmetric with stars of mass $1.23M_{\odot}$ and the APR4 EOS. From Kuan et al. 241.
The second remarkable feature of GRB 090510 is that of a rather prominent X-ray afterglow, as discussed by Rowlinson et al. 354, Suvorov and Kokkotas 396 and others. This afterglow exhibited the ‘internal plateau’ feature thought to be characteristic of a neutron star pumping spindown energy into the forward shock. Importantly, the system showed a steady decline in flux rather than a sharp cutoff, which is often interpreted as the collapse time of the metastable star [see, e.g. 343, 393]. If indeed a stable neutron star was formed, it is more plausible that the jet formation timescale was comparatively long, and hence that only one precursor was premerger with the second being attributable to a choking jet, possibly because neutron-stars have a harder time forming jets with high Lorentz factors [92]. Still, several scenarios are viable.
@@ V.6.6 GRB 230307A
GRB 230307A is an extremely bright, long GRB. The recorded gamma-ray fluence is $\gtrsim 3\times 10^{-3}\text{ erg cm}^{-2}$ in the 10–1000 keV band [119], second only to GRB 221009A. Despite its long duration, it is likely associated with a kilonova and heavy element production [253], thus resembling the case of GRB 211211A [119] of a merger-driven long GRB. Taking the best-fit distance of $\approx 291$ Mpc implies an enormous luminosity $L_{\rm prec}\sim 3.6\times 10^{50}\text{ erg s}^{-1}$. As discussed by Dichiara et al. [119], this event is a strong candidate for a magnetar with an external field of $\gtrsim 10^{15}$ G taking place in a merger. This is not only because of the spectrum and high luminoisty, but because of its long-short nature supporting a magnetar-in-a-merger origin if magnetic barriers are easier to set up (as for GRB 211211A). If the field was of this magnitude, accounting for magnetic shifting in mode eigenfrequencies would be essential, as would electromotive torques towards the end of the binary evolution (see Sec. IV.4). Given the waiting time of only $\approx 0.4$ s however, many ignition scenarios are viable (especially post-merger ones).
@@ V.7 Connections with post-merger phenomena
Insofar as strong magnetic fields are necessary to explain premerger, precursor spectra (Sec. V.4), the remnant environment and structure could be affected. It is fair to say that one of the open problems in GRB physics concerns whether or not magnetar remnants are capable of launching a relativistic jet that can drill through the polar baryon pollution [e.g. 359]. Most numerical simulations of mergers that leave a magnetar find that (i) a magnetic field of strength approaching equipartition ($\sim 10^{17}$ G) is needed to a launch a jet (if a jet can be launched at all), and (ii) the magnetic field that is generated is larger if the pre-merger ‘seed’ fields are larger [e.g. 91]. Point (ii) has been challenged by recent, high-resolution simulations [219], though it is naturally easier to accept strong remnant fields via flux conservation arguments if there is a magnetar in the merger.
The magnetic field strength of a remnant neutron-star would also influence its spindown energy and GW luminosity. In principle, one could anticipate a subpopulation of remnants with stronger fields if bright precursors are observed therefore [400]. Even if irrelevant for precursors, dynamo activity triggered by $f$- or other mode resonances just before merger could also amplify the seed field(s) (see Sec. VI.5).
@@ VI.1 Delay Timescales: postmerger jets
Fortunately — at least from a premerger perspective — there is reason to suspect that $t_{\rm J}$ may be considerably shorter than that seen in GRB 170817A in some instances, perhaps even down to a $\sim$ few ms. In this section we closely follow and summarise the compilations of Zhang 461 and Burns 73 in describing various jet formation mechanisms and $t_{\rm J}$ estimates. Where appropriate, some recent numerical simulations or theoretical results concerning remnants are instead quoted to update and/or adjust values presented in Table 1 of Zhang 461.
@@
Zhang 461 introduces a second decomposition for $\Delta t_{\rm jet}$, depending on the mechanism responsible for launching. The mechanism is naturally dependent on the progenitors. We write
@@
Table 4: Expected delays associated with merger-driven GRBs and relative precursor timings, adapted and expanded from Zhang 461. Abbreviations: NS = Neutron star; SM = Supramassive; HM = hypermassive. Remarks: (*) HMNS collapse time could theoretically be prolonged through thermal [313] or magnetic [393] support, extending the upper limit significantly. (**) Result quoted from analysis performed in Ciolfi et al. 92 regarding breakout requirements [see also Mösta et al. 293 and Ciolfi 90 for detailed discussions]. (†) Results anticipated from high-resolution dynamo simulations of Kiuchi et al. 219, who found amplification on sub-ms timescales. (‡) From simulations of Ruiz et al. 357, who found that sufficiently low-mass BHNSs result in a HMNS that tends to launch a magnetically-dominated jet. Argued by Shibata and Taniguchi 369 that for canonical neutron-star parameters, SGRB is limited to “low” energies of order $\sim 10^{48}$ erg/s because of tidal disruption impact on torus mass. (#) We assume a fireball-like mechanism for accretion scenarios involving a neutron star, or a Blandford-Znajek mechanism if involving a BH, and a Poynting-flux-dominated jet for dynamo (magnetic) scenarios when estimating $\Delta t_{\rm GRB}$. Though, in principle, any of these mechanisms could have an upper limit of many $\sim$ s owing to theoretical uncertainties; see Zhang 461 and Burns 73. Black hole topology and modified-gravity terms could also skew these estimates, though such effects are ignored [see, e.g. 376, 296].
@@
We close this section by noting that multimessenger events may accompany a collapse event for some of the remnant types listed in Tab. 4. Most notably, fast radio bursts (FRBs) have been predicted to be produced following the collapse of a neutron star [see, e.g. 139, 393]. Once an event horizon forms and cloaks the star, the external field lines are no longer anchored and so snap, inciting relativistic, magnetic shocks that accelerate electrons to high Lorentz factors, ultimately producing radiation in the $\gtrsim$ GHz band. A coincident detection of a sharp drop in X-ray flux and an FRB would be a “smoking gun” for a post-merger, neutron-star collapse; in such a case, at least $\Delta t_{\rm rem}$ could be tightly constrained. See Lu et al. 260 for recent (unsuccessful) search efforts for coincident FRBs and GRBs, though see Moroianu et al. 289 who found that FRB 20190425A is coincident with GW190425 at $\sim 2.8\sigma$. Sadly, no GRB was observed for this event.
@@ VI.2 Magnetospheric interaction and unipolar inductor
Depending on the persistence (or late-time generation) of strong (crustal) magnetic fields over cosmological timescales, electrodynamic interactions occurring in the final seconds of inspiral could produce fireworks. The mechanisms responsible have been described in detail by Hansen and Lyutikov 183 and others since. The production of gamma-rays can proceed through essentially three different channels, depending on the relative magnetization and alignment angles of the two stars [325, 244, 438]. In cases with a black hole, comparable estimates can be found in McWilliams and Levin 274.
The simplest, and arguably most probable if a luminous precursor is observed, scenario in the electrodynamic interaction context is that of the unipolar inductor introduced in Sec. IV.4. The reason for this is that it is already difficult to explain one magnetar in a merger let alone two, and without strong fields to mediate the interaction the precursor is likely to fizzle out long before it reaches Earth. The other two cases instead apply when the object’s have comparable magnetic dipole moments, either aligned or anti-aligned (though in reality of course there will be some angle, $\Theta_{\mu}$). In the aligned case, the field lines will compress at the interaction radius given by $r_{i}\approx a/(1+q_{\mu}^{1/3})$ with dipole-ratio $q_{\mu}=\mu_{B,d}/\mu_{A,d}$ [438]. In the opposite case, we have instead direct reconnection occurring in the interaction zone which leads to explosive event(s) followed by the establishment of a quasi-stable circuit similar to a unipolar inductor [e.g. 306, 99].
Lai 244 argues that the maximum energy that can be relieved relates to $\dot{E}_{\rm EM}$ from equation (47). This is controlled by the effective resistance between the two stars through the effective twist parameter $\zeta_{\phi}$, viz.
@@
where $B$ here is to be understood as the largest equatorial strength of either star and we have assumed equal stellar radii. The above is difficult to reconcile with precursor luminosities unless $a\sim 3R$ and $\zeta_{\phi}\gg 1$, though $\dot{E}_{\rm EM}$ can be slightly larger than estimates provided by Wang et al. 438 if $\zeta_{\phi}$ is kept free.
The estimate (55) is in rough agreement with that found from ab initio numerical simulations for large twists. Most and Philippov 291 found, via special-relativistic simulations, a maximum dissipation of (equation 6 therein)
@@
applying just $\sim$ ms before merger for equal mass objects [see also 99]. Note in particular the relevant $B$ value here is the local value at the interaction radius not that at the stellar surface (though these are approximately equal at merger by definition). Interestingly, the luminosity scaling was found by Most and Philippov 291 to be $\propto a^{-7/2}$, which is shallower than that of (55). Either way, we can match (56) with (55) at merger if $\zeta_{\phi}\lesssim 10$, consistent with the anti-aligned simulations by Palenzuela et al. 306, Crinquand et al. 99. Most and Philippov 292, Carrasco et al. 75 studied neutron-star plus black-hole binaries, finding roughly similar dissipation rates though with the interesting result that the flares present in their simulations were up to $\sim 10^{2}$ times brighter than the (relativistic) orbital emission.
An important aspect of the interaction model is that emissions should be generated almost instantaneously. That is, particle production occurs immediately after reconnection and thus the precursor timing should correspond directly to the ignition moment without an additional lag beyond those detailed in Sec. VI.1. On the other hand, the duration of a precursor born from magnetospheric interaction is likely limited to $\sim 10$ ms based on the relationship between the Poynting flux and chirp length where the separations are low [142]. This is somewhat shorter than all event durations listed in Tab. 3, though allowing for a factor $\sim$ few extension would bring many low-waiting-time precursors into view (e.g. 150922A, 100223110 , 100702A, …). Still, Metzger and Zivancev 283 argue the detection horizon is limited to $\sim 10$ Mpc for $B\approx 10^{14}$ G unless the twist is large, $\zeta_{\phi}\gg 1$.
Finally, aside from reconnection being generally able to spark a plethora of progenitors (see also Sec. VII), strong magnetic fields in the emission environment could effect photon propagation directly. For example, quantum-mechanical photon splitting could acquire a non-negligible cross section [108, 275], thereby impacting on the polarization states and/or Faraday rotation of electromagnetic waves. Such effects, manifesting as vacuum birefringence, have been observed in emissions from the isolated neutron stars RX J1856.5-3754 [286] and 4U 0142+61 [410] with characteristic field strengths of a few by $10^{13}$ G [330] and $\sim 10^{14}$ G [304], respectively. As noted by Wang and Liu 436, the next-generation gamma-ray polarimeter POLAR-2 may be able to measure such effects directly and constrain the plasma environment [202].